OpenStudy (anonymous):

3 − 5(x + 2) = 3(x + 1) + 6 explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

you know the answer stop playing!..lol

OpenStudy (anonymous):

ik just dont know how to explain it 4 my test

OpenStudy (anonymous):

lol Thanks for giving me a medal for no reason!..lol :)

OpenStudy (anonymous):

how did you solve the equation?

OpenStudy (anonymous):

np and you would have to -3 frm both sides right

OpenStudy (anonymous):

right

OpenStudy (anonymous):

next

OpenStudy (anonymous):

use those big brains you have!..lol :)

OpenStudy (anonymous):

hold on brb

OpenStudy (anonymous):

ok then 5x+10=3x+3+6 5x+10=3x+9 5x=3x-1 2x=-1 x-2

OpenStudy (anonymous):

you got it!